The viscosity of a fluid is to be measured by a viscometer constructed of two 3m long concentric cylinders.the inner diameter of the outer cylinder is 35cm, and the gap between the two cylinders is 0.14cm.the outer cylinder is rotated at 250rpm. and the torque is measured to be 1.8N determine the viscosity of the fluid

1
Expert's answer
2022-04-13T09:21:09-0400

"\eta=\frac {2Ms}{l\pi d^2\omega r}=34~\frac{m^2}s."